# Universal Entanglement Revival of Topological Origin

### v1

@@ Abstract
We have analyzed the dynamics of entanglement in dissipative fermionic and bosonic Su–Schrieffer–Heeger (SSH) models and found that, when the decoherence channel preserves the chiral symmetry, they exhibit a revival of entanglement in a wide range of parameters. This behavior only emerges in the topological phase, with the visibility of the revival dropping to zero at the phase boundary. Furthermore, the revival acquires a universal character once the system size exceeds the localization length of the edge modes. Our findings indicate that the universal entanglement revival has its origin in the topological properties of the SSH model. These dynamical properties may be experimentally accessible, for example, using photonic quantum computers.
1 Introduction
Entanglement is a critical resource in quantum information processing, yet is intrinsically fragile with respect to interactions with the environment. These induce decoherence, which typically leads to the rapid degradation of entanglement. Nevertheless, certain systems have been shown to exhibit the interesting phenomenon of entanglement revival, in which, after the initial decay, an unexpected recovery occurs at a later stage of the dynamical evolution [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. A robust mechanism causing this type of dynamics has been recently discussed for an ensemble of three-level systems interacting with single-mode bosons [16], where the revival is associated to the presence of a special zero-energy state weakly affected by decoherence. The dissipative evolution rapidly destroys the initial entanglement but eventually turns itself into a self-purification process, resulting in a distinct entanglement revival.
The mechanism of Ref. [16] is generic, thus might be applicable to other scenarios. Indeed, we show here that topological system are a natural candidate to observe this type of revivals. The potential relevance of Ref. [16] for topological systems is suggested by the crucial role played in the three-level ensemble by a chiral symmetry, and by the fact that the state responsible for the revival belongs to the zero-energy subspace associated with that symmetry. In this sense, the metastable state of Ref. [16] is analogous to the zero-energy edge mode of a topological system.
To explore concretely these ideas, we specialize our analysis to a dissipative version of the Su-Schrieffer-Heeger (SSH) model. The SSH model is a well-known paradigm for topological insulators [17, 18, 19] and has considerable relevance for quantum information and quantum computing applications [20, 21, 22, 23, 24, 25], due to its robustness against disorder and defects [26, 27, 28, 29, 30, 31, 32]. Recent studies have increasingly focused on how topological systems, including the SSH model, behave under dissipative dynamics [33, 34, 35, 36, 37, 38]. Gaining a deeper understanding of dissipation in topologically protected systems is critical for furthering the practical implementation of quantum technologies. Here we explore how the topological properties of the SSH model influence entanglement dynamics, which is of significant interest for the development of more resilient quantum systems.
@@ 1 Introduction
The topological edge states of the SSH model remain stable when decoherence only acts on one of the two sublattices (specifically, the odd sites). Under this scenario, we find the occurrence of an entanglement revival directly tied to the presence of a topological edge mode. We also show that the revival approaches a universal form when increasing the system size, i.e., the evolution of entanglement becomes independent of the chain length. Therefore, the presence (absence) of the revival is a definite signature of the topological (trivial) phase. Our findings apply to both fermionic and bosonic versions of the SSH model, thus they could be simulated in a controlled manner using photonic quantum computers [39, 40, 41]. Nearest-neighbor hoppings of variable coupling strengths can be realized with tunable beam splitters. Additionally, the required photon loss pattern — lower at even sites and higher at odd sites — can be simulated by selective leakage of photons. Therefore, we expect that the topologically protected entanglement revivals discussed here are achievable with current photonic technology.
@@
This paper is organized as follows: In Section 2, we introduce the SSH model, briefly reviewing its chiral symmetry and topological properties. The theoretical methods we used are detailed in Section 3, including the semiclassical approach and different entanglement measures. Section 4 presents the results, highlighting the universal entanglement revival observed in both fermionic and bosonic SSH models as well as its connection to topological edge states. Section 5 concludes the paper. To further support the arguments in the main text, two appendices are provided: A demonstrates the universal revival behavior in the mutual information as an alternative extent of entanglement. B examines the density of von Neumann and Rényi entropies, which also exhibit a universal behavior.
2 Model
@@
Figure 1: Schematics of the dissipative Su-Schrieffer-Heeger (SSH) model. The coupling strengths between nearest-neighbor sites alternate between two values, $g_{1}$ and $g_{2}$, and the quantum decoherence process (wavy lines) occurs only at the odd sites. For numerical efficiency, we ignore the shaded region and denote the remaining part as ‘half-SSH chain’.
@@
The SSH model is a prototype of one-dimensional (1D) topological insulators [18, 19]. As a tight-binding model with alternating hopping amplitudes, $g_{1}$ and $g_{2}$, it exhibits a chiral symmetry (also known as sub-lattice symmetry), which, together with the time-reversal symmetry, completely classifies the topological states of matter. The chiral symmetry is characterized by a unitary operator that anti-commutes with the Hamiltonian, and ensures that the Hamiltonian matrix can be written in a block-off-diagonal form, leading to a degenerate zero-energy subspace and a symmetric energy spectrum around zero energy. The zero-energy states are especially interesting: According to the bulk-boundary correspondence principle, they always accompany the topologically non-trivial phase of topological insulators and are localized at the boundaries of the system. In the SSH model, the topological phase occurs when $g_{1}<g_{2}$  [^1] Without loss of generality, we assume real and positive hopping amplitudes. Complex hopping amplitudes can always be brought to real positive values with a gauge transformation.. In this phase, there are two zero-energy edge states at the two ends of the chain. The localization length of each edge state describes how far the state spreads from the edge into the bulk of the chain, and is given by $\xi=2/\log[g_{2}/g_{1}]+1$.
@@
For systems of sufficiently large size $L\gg\xi$, the zero-energy edge states at the two ends are decoupled from each other. As the effects of interest in this work only involve a single edge state, we will, for the sake of numerical efficiency, consider only half of the physically realistic chain, ignoring the shaded portion in Fig. 1  [^2] In this half-chain model one unit cell is not complete, as it contains only a single site instead of two. Therefore, the bulk-boundary correspondence does not hold in the usual sense.. This simplification allows us to concentrate on the properties and behavior of a single edge state, while maintaining the essential features of the topological phase. This half-SSH model contains $n$ two-site unit cells and a single unpaired site, so the overall chain length $L=2n+1$ is always odd. The Hamiltonian reads:
@@
where $\hat{f}_{i}^{\dagger}$ and $\hat{f}_{i}$ are, respectively, the creation and annihilation operators at site $i=0,1,\ldots,L-1$. In the fermionic (bosonic) SSH model, they satisfy the anti-commutation (commutation) relations, $\{\hat{f}_{i},\hat{f}_{j}^{\dagger}\}=\delta_{ij}$ $([\hat{f}_{i},\hat{f}_{j}^{\dagger}]=\delta_{ij}).$ To better reveal the consequences of the chiral symmetry and the topological properties of the system, we rewrite the Hamiltonian in terms of the normal modes $\hat{C}_{\pm k}$:
@@
where the single-particle energies ±$\pm\epsilon_{k}$ are given by:
@@
with $\tilde{k}:=2\pi k/(L+1).$ The normal modes are related to the site operators $\hat{f}_{i}$ by
@@
and $\cos(\phi_{k}):=(g_{1}\cos\tilde{k}+g_{2})/\epsilon_{k}.$ Furthermore, there is a special normal mode with zero energy, which does not appear in Eq. (2):
@@
Clearly, when $g_{1}<g_{2}$ this mode has its probability amplitudes localized at the left end of the system  [^3] In the full SSH model, there is another zero-energy edge mode at the right end.. For later use, when $k=0$ we define:
@@
Recall that, unlike this zero-energy edge mode, all other modes $\hat{C}_{\pm k}$ ($k\neq 0$) are bulk states with probability amplitudes spread over the whole chain.
We are interested in the topological influence on the dynamics of quantum entanglement in the presence of decoherence. To clearly illustrate the topological effects, we assume that the quantum decoherence processes preserve the chiral symmetry; that is, we assume that decoherence processes occur only on the odd sites $\hat{f}_{2i-1}$, with a uniform decay rate $\Gamma$. The noisy dynamics is governed by the quantum master equation
@@
where $\hat{\rho}$ is the density operator of the system and $\mathcal{D}$ is the superoperator $\mathcal{D}[\hat{f}]\hat{\rho}=\hat{f}\hat{\rho}\hat{f}^{\dagger}-(\hat{f}^{\dagger}\hat{f}\hat{\rho}+\hat{\rho}\hat{f}^{\dagger}\hat{f})/2,$ describing the quantum decoherence processes.
We will consider an initial state of the form:
$|\psi_{g}\rangle=\hat{C}_{0}^{\dagger}\Pi_{k=1}^{(L-1)/2}\hat{C}_{-k}^{\dagger}|\rangle,$ (9)
where |$\rangle$| ⟩ is the vacuum state. Since all negative-energy modes and the zero-energy mode are singly occupied, this initial state is one of the two degenerate many-body ground states of the fermionic model. In that case, it is a natural choice for the initial state. In the bosonic model, occupying all negative and zero-energy normal modes can be achieved by using external driving, as commonly done in experimental demonstrations of topological band structures, e.g., in photonic crystals [42, 43]. Since we study non-equilibrium dynamics, the resulting physical effects inevitably depend on the initial state. However, as long as the zero-energy edge mode is initially occupied, our main findings remain qualitatively valid for many other initial states. In particular, we have checked that a universal revival is found for an initial state where all the normal modes are singly occupied. This confirms that the universal entanglement dynamics reported below is closely related to topological features.
3 methods
Before presenting in Sec. 4 our main results, we describe in this Section the methods used to solve the quantum master equation, Eq. (8), and to analyze the entanglement content of the solution.
3.1 Semiclassical approach
The computational cost to exactly solve the quantum master equation increases exponentially with system size  [^4] For non-interacting fermions with a restricted class of quantum jump operators, a polynomial-time method is available. See, for example, Refs. [44, 45].. Fortunately, in the parameter regime of our interest, a semicassical approach provides an accurate description of the dynamics. We first note that, when the coupling strengths are much larger than the decay rate ($g_{1},g_{2}\gg\Gamma$), the eigenstates of Eq. (1) are generally well-separated in energy. Since the coherent phase factor between two eigenstates oscillates rapidly in time, it gets averaged out throughout the slow dissipative evolution and we may assume that $\hat{\rho}(t)$ is approximately diagonal in the many-body eigenstates, $\mathinner{\lvert{\textstyle E_{\alpha}}\rangle}$:
@@ 3.1 Semiclassical approach
The probabilities $P_{\alpha}$ of different many-body eigenstates $|E_{\alpha}\rangle$ satisfy the master equation
@@
where the transition rate $\gamma_{\alpha\beta}$ from $\mathinner{\lvert{\textstyle E_{\beta}}\rangle}$ to $\mathinner{\lvert{\textstyle E_{\alpha}}\rangle}$ is given by $\gamma_{\alpha\beta}:=\Gamma\sum_{i=1}^{(L-1)/2}|\langle E_{\alpha}|\hat{f}_{2i-1}|E_{\beta}\rangle|^{2}.$ In terms of the eigenmode operators:
$\gamma_{\alpha\beta}:=\frac{\Gamma}{2}\sum_{k\neq 0}|\langle E_{\alpha}|\hat{C}_{k}|E_{\beta}\rangle|^{2},$ (12)
which can be derived using $\sum_{i=1}^{(L-1)/2}B_{ki}B_{k^{\prime}i}=\delta_{kk^{\prime}}/2.$ This relation has its origin in the the chiral symmetry, leading to coefficients $A_{ki}$ and $B_{ki}$, see Eqs. (5) and (7), which are separately normalized. Therefore, we also have $\sum_{i=0}^{(L-1)/2}A_{ki}A_{k^{\prime}i}=\delta_{kk^{\prime}}/2.$
We note some unique properties of the semiclassical dynamics described by Eqs. (10) and (11). First, every time a decay process occurs, the total number of particles decreases by one. However, the zero-energy edge state can never decay, because $\hat{C}_{0}$ has non-zero amplitudes only at the even sites, see Eq. (6). In the long-time limit ($t\to\infty$) only one particle remains, occupying the zero-energy edge state:
$\hat{\rho}_{\infty}:=\lim_{t\rightarrow\infty}\hat{\rho}(t)=\hat{C}_{0}^{\dagger}|\rangle\langle|\hat{C}_{0}.$ (13)
Second, with the initial condition Eq. (9) the occupation numbers of the eigenmodes are either one or zero, and a decay process corresponds to one of the occupied modes becoming empty. From Eq. (12), we see that all these decays occur with an identical rate, $\Gamma/2$, which allows us to solve the semiclassical evolution of Eqs. (10) and (11) in closed form:
$\hat{\rho}(t)\approx\sum_{m=0}^{(L-1)/2}P_{m+1}(t)\sum_{\{k_{i}\}_{i=1}^{m}}\hat{C}_{\{k_{i}\}}^{\dagger}\hat{\rho}_{\infty}\hat{C}_{\{k_{i}\}},$ (14)
@@
$P_{m+1}(t)=\exp\left[\frac{-m\Gamma t}{2}\right]\left(1-\exp\left[-\frac{\Gamma t}{2}\right]\right)^{\frac{L-1}{2}-m}.$ (15)
In Eq. (14), we used the notation $\hat{C}_{\{k_{i}\}_{i=1}^{m}}:=\prod_{i=1}^{m}\hat{C}_{-k_{i}}$ where $\{k_{i}\}_{i=1}^{m}$ is an unordered set of indexes (with $k_{i}>0$). We see from Eq. (14) that the probabilities $P_{\alpha}$ of the eigenstates hosting $m-1$ particles with negative energy (and one particle in the $\hat{C}_{0}$ mode) are all equal. For those states we can simply set $P_{\alpha}=P_{m}$.
So far, we have justified the semiclassical approximation in the large-coupling limit $g_{1},g_{2}\gg\Gamma$. However, the above description is more general, and allows us to compute the dynamics deep in the topological phase ($g_{1}\ll g_{2}$), when $g_{1}$ can become comparable to $\Gamma$. In this limit, the energy differences between the eigenstates with the same $m$ appearing in Eq. (14) are of order $g_{1}$, thus the corresponding coherences would not oscillate on a fast timescale. Fortunately, however, the dissipative evolution does not generate such coherences, as can be checked by applying the full master equation, Eq. (8), to a generic eigenstate $|E_{\alpha}\rangle\langle E_{\alpha}|$. This is due to a destructive cancellation of amplitudes, which follows from the orthogonality of the eigenmodes. In particular, $\sum_{i=1}^{(L-1)/2}B_{ki}B_{k^{\prime}i}=0$ when $k\neq k^{\prime}$. The semiclassical approximation is still not exact, as the full dynamics generates some coherence between eigenstates with large energy differences, of order $g_{2}$. Thus, it is still necessary to satisfy the condition $g_{2}\gg\Gamma$.
To demonstrate the accuracy of the semiclassical approximation when $g_{1}\approx\Gamma$, we compare in Fig. 2 two results for the probability $P_{m}(t)$: One from the numerical solution of the full quantum master equation (8) and another from the semiclassical approximation (15). Clearly we observe an excellent agreement, justifying the semiclassical approach. In passing, we note that the maximum probability appears earlier for states with more particles, which is intuitively appealing.
@@
The semiclassical state in Eq. (14) provides a complete description of the dynamics. For fermionic systems, the state satisfies Wick’s theorem [46], hence one can further rewrite it into a Gaussian form [47]:
@@
[…]
$\chi_{k}(t)=\begin{cases}1-\exp(-\Gamma t/2)&{\rm for}\leavevmode\nobreak\ k<0,\\ 0&{\rm for}\leavevmode\nobreak\ k=0,\\ 1&{\rm for}\leavevmode\nobreak\ k>0.\end{cases}$ (18)
@@
The Gaussian form (16) of the density matrix turns out to be extremely powerful when later calculating the entanglement content. We emphasize that for bosons the mixed state in Eq. (14) is not Gaussian anymore, despite having exactly the same form.
3.2 Entanglement measures

### v2

@@ Abstract
We investigate the dynamics of entanglement in dissipative fermionic and bosonic Su–Schrieffer–Heeger (SSH) models and discover that they exhibit a revival dynamics of entanglement when the decoherence channel preserves the chiral symmetry. This behavior is only observable in the topological phase, and the visibility of the revival diminishes to zero at the phase boundary. Furthermore, the revival acquires a universal character, meaning that the entire time-evolution profile remains independent of system size, provided that the size exceeds the localization length of the edge modes associated with the topological phase. Our findings suggest that the universal entanglement revival originates from the topological properties of the SSH model. These dynamical properties may be experimentally accessible, for instance, by utilizing photonic quantum computers.
I Introduction
Quantum entanglement, a critical resource in quantum science and technology, is fragile and prone to rapid degradation under quantum decoherence Zurek (2003); Schlosshauer (2007). Consequently, systems exhibiting entanglement revival phenomena, where quantum correlations reemerge following initial decay despite persistent decoherence, have attracted significant interest Życzkowski et al. (2001); Tanaś and Ficek (2004). Ever since, this revival behavior has been extensively investigated across diverse quantum platforms, spanning both discrete- and continuous-variable systems Benatti and Floreanini (2006); Ficek and Tanaś (2006); Cunha (2007); Paz and Roncaglia (2008); Drumond and Cunha (2009); Das and Agarwal (2009); Orszag and Hernandez (2010); Amico et al. (2018); Aolita et al. (2015). Studies have demonstrated even more striking dynamical features, including entanglement sudden death and rebirth phenomena Ficek and Tanaś (2008); Abdel-Aty and Yu (2008); López et al. (2008); Aolita et al. (2015); Lakhfif et al. (2022); Nunavath et al. (2023). Such entanglement revivals must be distinguished from the direct influence of interaction which, by its very nature, also tends to enhance entanglement. The latter acts on a time scale given by the inverse interaction strength, whereas the entanglement revivals happen over a much longer time scale (typically, longer than the quantum decoherence time).
Various physical mechanisms have been suggested to understand the occurrence of entanglement revival behavior, which usually involves a non-trivial interplay between coherent interactions and environmental decoherence Życzkowski et al. (2001); Tanaś and Ficek (2004); Benatti and Floreanini (2006); Ficek and Tanaś (2006); Cunha (2007); Paz and Roncaglia (2008); Drumond and Cunha (2009); Das and Agarwal (2009); Orszag and Hernandez (2010); Amico et al. (2018); Aolita et al. (2015); Ficek and Tanaś (2008); Abdel-Aty and Yu (2008); López et al. (2008); Lakhfif et al. (2022); Nunavath et al. (2023); Chen et al. (2024). Of our particular interest is a generic scenario based on the chiral symmetry, recently discussed for an ensemble of three-level systems interacting with single-mode bosons Chen et al. (2024). In this scenario, the chiral symmetry generates a distinct zero-energy state that demonstrates remarkable resilience against decoherence. Initially, the system experiences a phase where dissipation rapidly diminishes the initial entanglement. Subsequently, the system undergoes a gradual transition to the chiral symmetry-protected zero-energy state, eventually undergoing a self-purification process that entails a distinct revival of entanglement.
Chiral symmetry plays an especially prominent role in symmetry-protected topological states of matter Chiu et al. (2016). In this context, the associated zero-energy states become localized at the system’s edges and, according to the bulk-boundary correspondence principle, serve as distinctive signatures of topological phases. Entanglement is both a fundamental quantum resource and a valuable tool for characterizing quantum phase transitions Osterlo et al. (2002); Amico et al. (2008) as well as topological phase transitions Kitaev and Preskill (2006); Fendley et al. (2007). Therefore, by combining these perspectives, it is natural to expect that the entanglement revival dynamics can provide valuable insights into the characterization of topological phases protected by chiral symmetry.
Motivated by this idea, in the present work, we investigate entanglement dynamics in a dissipative variant of the Su-Schrieffer-Heeger (SSH) model. The SSH model represents one of the best-known examples of chiral symmetry-protected topological insulators Su et al. (1979); Asbóth et al. (2016); Shen (2017), and has considerable relevance for quantum technology applications Kitaev (2003); Michael H. Freedman and Wang (2003); Nayak et al. (2008); Bomantara and Gong (2018a, b); Lahtinen and Pachos (2017), due to its simple form and robustness against disorder Halperin (1982); Büttiker (1988); Niu and Thouless (1984); Hasan and Kane (2010); Qi and Zhang (2011); Ryu et al. (2010); Kawasaki et al. (2022). Recent studies have extended topological analysis to dissipative dynamics Huang et al. (2020); Diehl et al. (2011); Bardyn et al. (2013); Rivas et al. (2013); Viyuela et al. (2012); Dangel et al. (2018); Salatino et al. (2025), yet the interplay between topological properties and entanglement dynamics remains underexplored. Understanding this relationship is critical for implementing topological systems in practical quantum technologies. Our work systematically explores how the topological phases manifest through entanglement revival behaviors in the SSH model, establishing a characterization of topological phases in terms of entanglement dynamics.
We find that, in the topological phase of the SSH model, the dissipative entanglement dynamics not only exhibits a revival behavior but also takes a universal form; in other words, the entire time-evolution profile of entanglement is independent of the chain length, with curves for different system sizes collapsing onto a single curve. The universal profile becomes evident for chain lengths exceeding the localization length of the zero-energy edge modes, which are characteristic of the topological phase according to the bulk-boundary correspondence Chiu et al. (2016). In the trivial phase, however, all these features disappear. All these suggest that the entanglement revival behavior is directly tied to the presence of the zero-energy edge modes, i.e., the topological state of the SSH model.
@@ I Introduction
Our findings apply to both fermionic and bosonic versions of the SSH model, thus they can be simulated in tunable quantum systems, such as photonic quantum computers Madsen et al. (2022); O’Brien et al. (2009); Yuan et al. (2024). On a photonic quantum computer, nearest-neighbor hoppings of variable coupling strengths can be realized with tunable beam splitters. Additionally, the required photon loss pattern with lower and higher rates at even and odd sites, respectively, can be simulated by selective leakage of photons. Therefore, we expect that the topologically protected entanglement revival reported in this work is achievable with current photonic technology. We also note that a fermionic SSH model with tunable couplings has been experimentally realized using donor-based quantum-dot arrays in silicon Kiczynski et al. (2022), making this another promising platform for testing our predictions.
@@
This paper is organized as follows: In Section II, we introduce the SSH model, briefly reviewing its chiral symmetry and topological properties. The theoretical methods we used are detailed in Section III, including the semiclassical approach and different entanglement measures. Section IV presents the main results, highlighting the universal entanglement revival observed in both fermionic and bosonic SSH models, as well as its connection to topological edge states. Section V concludes the paper. To further support the arguments in the main text, four appendices are provided: Appendix A demonstrates the universal revival behavior in the mutual information, as an alternative to entanglement. Appendix B examines the von Neumann and Rényi entropies per site (i.e., the entropy densities), which also exhibit a universal behavior. Appendices C and D investigate more realistic conditions, including spatially correlated noise and finite temperature, and confirm that the universal revival remains qualitatively robust.
II Model
@@
Figure 1: Schematics of the dissipative Su-Schrieffer-Heeger (SSH) model. The coupling strengths between nearest-neighbor sites alternate between two values, $g_{1}$ and $g_{2}$, and the quantum decoherence process (wavy lines) occurs only at the odd sites. For numerical efficiency, we ignore the shaded region and refer to the remaining part as the ‘half-chain SSH mdoel’.
@@
The SSH model is a prototype of one-dimensional (1D) topological insulators Asbóth et al. (2016); Shen (2017). As a tight-binding model with alternating hopping amplitudes, $g_{1}$ and $g_{2}$, it exhibits a chiral symmetry (also known as sub-lattice symmetry), which, together with the time-reversal symmetry, completely classifies the topological states of matter. The chiral symmetry is characterized by a unitary operator that anti-commutes with the Hamiltonian, and ensures that the Hamiltonian matrix can be written in a block-off-diagonal form, leading to a degenerate zero-energy subspace and a symmetric energy spectrum around the zero energy. The zero-energy modes are especially interesting: According to the bulk-boundary correspondence principle, they always accompany the topologically non-trivial phase of topological insulators and are localized at the boundaries of the system. In the SSH model, the topological phase occurs when $g_{1}<g_{2}$  [^1] Without loss of generality, we assume real and positive hopping amplitudes. Complex hopping amplitudes can always be brought to real positive values with a gauge transformation.. In this phase, there are two zero-energy edge modes at the two ends of the chain. The localization length of each edge mode,
$\xi=1+\frac{2}{\log(g_{2}/g_{1})},$ (1)
@@
describes how far the mode spreads from the edge into the bulk of the chain. For systems of sufficiently large size $L\gg\xi$, the zero-energy edge modes at the two ends are decoupled from each other. In this case, as long as any physical effects due to the zero-energy edge modes are concerned, one can focus on each edge mode separately. Therefore, for the sake of numerical efficiency, we will consider only half of the physically realistic chain, ignoring the shaded portion in Fig. 1  [^2] In this half-chain SSH model, one of the unit cells is not complete, as it contains only a single site instead of two, and the bulk-boundary correspondence does not hold in the usual sense.. This simplification allows us to concentrate on the properties and behavior of a single edge mode, while maintaining the essential features of the topological phase. This half-chain SSH model contains $n$ two-site unit cells and a single unpaired site, so the overall chain length $L=2n+1$ is always odd. The Hamiltonian reads as
@@
where $\hat{f}_{k}^{\dagger}$ and $\hat{f}_{i}$ are the creation and annihilation operators, respectively, at site $i=0,1,\ldots,L-1$. In the fermionic (bosonic) SSH model, they satisfy the anti-commutation (commutation) relations, $\{\hat{f}_{i},\hat{f}_{j}^{\dagger}\}=\delta_{ij}$ $([\hat{f}_{i},\hat{f}_{j}^{\dagger}]=\delta_{ij}).$ To better reveal the consequences of the chiral symmetry and the accompanying topological properties of the system, we rewrite the Hamiltonian in terms of the normal modes $\hat{c}_{\pm k}$:
@@
where the single-particle energies $\pm\epsilon_{k}$ are given by:
@@
with $\tilde{k}:=2\pi k/(L+1).$ The normal modes are related to the bare fermion/boson modes $\hat{f}_{i}$ by
@@
and $\cos(\phi_{k}):=(g_{1}\cos\tilde{k}+g_{2})/\epsilon_{k}.$ Furthermore, there is a special normal mode which does not appear in Eq. (3) as it has the zero energy:
@@
Clearly, when $g_{1}<g_{2}$, this zero-energy mode has its probability amplitudes localized at the left end of the system  [^3] In the full SSH model, there is another zero-energy edge mode at the right end.. For later use, when $k=0$ we define:
@@
Recall that, unlike this zero-energy edge mode, all other modes $\hat{c}_{\pm k}$ ($k\neq 0$) are bulk modes with probability amplitudes spread over the whole chain.
We are interested in the topological influence on the dynamics of quantum entanglement in the presence of decoherence. To clearly illustrate the topological effects, we assume that the quantum decoherence processes preserve the chiral symmetry; that is, we assume that decoherence processes occur only on the odd sites, 2i1{2i-1}, with a uniform decay rate $\Gamma$. The noisy dynamics is governed by the quantum master equation for the density operator $\hat{\rho}$
@@
where $\mathcal{D}$ is the superoperator
[…]
describing the single-particle decay process and $\Gamma$ is the decay rate.
We will consider an many-body initial state of the form:
$|\psi_{g}\rangle=\hat{c}_{0}^{\dagger}\prod_{k=1}^{(L-1)/2}\hat{c}_{-k}^{\dagger}\,\mathinner{\lvert{\textstyle\ }\rangle}$, (11)
[…]
III Methods
Before presenting in Sec. IV our main results, here we describe the methods used to solve the quantum master equation, Eq. (9), and to analyze the entanglement content of the solution.
III.1 Semiclassical approach
The computational cost to exactly solve the quantum master equation increases exponentially with system size  [^4] For non-interacting fermions with a restricted class of quantum jump operators, a polynomial-time method is available; See, for example, Refs. Prosen (2008); Bravyi and Knönig (2012); Alba and Carollo (2023).. Fortunately, in the parameter regime of our interest, a semicassical approach provides an accurate description of the dynamics. We first note that, when the coupling strengths are much larger than the decay rate ($g_{1},g_{2}\gg\Gamma$), the eigenstates of Eq. (2) are generically well-separated in energy. Since the coherent phase factor between energy-separated eigenstates oscillates rapidly in time, it gets averaged out throughout the slow dissipative evolution. We can therefore assume that $\hat{\rho}(t)$ is approximately diagonal in the many-body eigenstates, $\mathinner{\lvert{\textstyle E_{\alpha}}\rangle}$:
@@ III.1 Semiclassical approach
The probabilities $P_{\alpha}$ for different many-body eigenstates $|E_{\alpha}\rangle$ satisfy the master equation
@@
where the transition rate $\gamma_{\alpha\beta}$ from $\mathinner{\lvert{\textstyle E_{\beta}}\rangle}$ to $\mathinner{\lvert{\textstyle E_{\alpha}}\rangle}$ is given by
$\gamma_{\alpha\beta}:=\Gamma\sum_{i=1}^{(L-1)/2}|\langle E_{\alpha}|\hat{f}_{2i-1}|E_{\beta}\rangle|^{2}.$ (14)
In terms of the normal-mode operators $\hat{c}_{k}$, it reads as
$\gamma_{\alpha\beta}=\frac{\Gamma}{2}\sum_{k\neq 0}|\langle E_{\alpha}|\hat{c}_{k}|E_{\beta}\rangle|^{2},$ (15)
which follows from the orthogonality of coefficients $B_{ki}$,
$\sum_{i=1}^{(L-1)/2}B_{ki}B_{k^{\prime}i}=\delta_{kk^{\prime}}/2.$ (16)
In fact, this orthogonality relation has its origin in the chiral symmetry, leading to the separately normalized coefficients $A_{ki}$ and $B_{ki}$; see Eqs. (6) and (8). Therefore, we also have
$\sum_{i=0}^{(L-1)/2}A_{ki}A_{k^{\prime}i}=\delta_{kk^{\prime}}/2.$ (17)
We note some unique properties of the semiclassical dynamics described by Eqs. (12) and (13): First, every time a decay process occurs, the total number of particles decreases by one. However, the zero-energy edge mode can never decay, because $\hat{c}_{0}$ has non-zero amplitudes only at the even sites, see Eq. (7). In the long-time limit ($t\to\infty$), only one particle remains, occupying the zero-energy edge mode; i.e.,
$\hat{\rho}_{\infty}:=\lim_{t\rightarrow\infty}\hat{\rho}(t)=\hat{c}_{0}^{\dagger}\mathinner{\lvert{\textstyle\ }\rangle}\mathinner{\langle{\textstyle\ }\rvert}\hat{c}_{0}.$ (18)
Second, with the initial condition in Eq. (11), the occupation numbers of the normal modes are either one or zero, and a decay process corresponds to one of the occupied modes becoming empty. From Eq. (15), we see that all these decays occur with an identical rate, $\Gamma/2$, which allows us to solve the semiclassical evolution of Eqs. (12) and (13) in a closed form as
[…]
@@
$P_{m+1}(t)=e^{-{m\Gamma t}/{2}}\left(1-e^{-{\Gamma t}/{2}}\right)^{\frac{L-1}{2}-m}.$ (20)
In Eq. (19), we have used the short-hand notation
[…]
where $\{k_{i}\}_{i=1}^{m}$ is an unordered set of mode indices (with $k_{i}>0$). We see from Eq. (19) that the probabilities $P_{\alpha}$ for the many-body eigenstates $\alpha$ to host m1m-1 particles at negative energies (and one particle in the $\hat{c}_{0}$ mode) are all equal. For those states, we can simply set $P_{\alpha}=P_{m}$.
So far, we have justified the semiclassical approximation in the large-coupling limit $g_{1},g_{2}\gg\Gamma$. However, the above description is more general, and allows us to compute the dynamics deep in the topological phase ($g_{1}\ll g_{2}$), when $g_{1}$ can become comparable to $\Gamma$. In this limit, the energy differences between the eigenstates with the same $m$ appearing in Eq. (19) are of order of $g_{1}$, thus the corresponding coherent phase factors would not oscillate on a fast timescale. Fortunately, however, the dissipative evolution does not generate such coherence, as can be checked by applying the full master equation, Eq. (9), to a generic eigenstate $|E_{\alpha}\rangle\langle E_{\alpha}|$. This is due to a destructive interference of amplitudes, which follows from the orthogonality of the normal modes; i.e., Eqs. (16) and (17). Nevertheless, it is still necessary to satisfy the condition $g_{2}\gg\Gamma$ in order to prevent the full dynamics from generating coherence between many-body eigenstates with large energy differences of order $g_{2}$.
To demonstrate the accuracy of the semiclassical approximation when $g_{1}\approx\Gamma$ (yet $g_{2}\gg\Gamma$), we compare in Fig. 2 the numerical solution of the full quantum master equation (9) and the semiclassical approximation (20). Clearly, we observe an excellent agreement between the two, justifying the semiclassical approach. The validity of the semiclassical approximation can be understood in terms of two complementary mechanisms. First we note that the coherence between eigenstates with the same number of particles (occupying the normal modes $\hat{c}_{i}$) vanishes due to destructive interference between decay paths into different eigenstates. This property can be derived from Eqs. (16) and (17). Specifically, when the system evolves from one initial state into a coherent superposition of such eigenstates, the amplitudes associated with different paths acquire opposite phases. These phase differences, which originate from the symmetry and spatial structure of the modes, lead to a cancellation of coherence. Furthermore, for eigenstates with different particle-numbers, coherence terms are suppressed by rapid oscillations arising from large energy gaps, set primarily by $g_{2}$. As long as $g_{2}\gg\Gamma$, these coherence terms effectively average out. In passing, we note that the maximum probability appears earlier for initial states with more particles, which is intuitively appealing.
@@
The semiclassical state in Eq. (19) provides a complete description of the dynamics. For fermionic systems, the state satisfies Wick’s theorem Mahan (2000), hence one can further rewrite it into a Gaussian form Peschel (2003):
@@
Here, $G$ denotes the matrix of single-particle Green’s functions,
$G_{ij}(t):=\left\langle\hat{f}_{i}\hat{f}_{j}^{\dagger}\right\rangle=\operatorname{{Tr}}\left[\hat{\rho}(t)\hat{f}_{i}\hat{f}_{j}^{\dagger}\right],$ (24)
which can be conveniently obtained by first evaluating the Green’s functions of the normal modes $\hat{c}_{k}$. Since the density matrix $\hat{\rho}$ in Eq. (19) is diagonal in the normal modes, we immediately find that $\chi_{kk^{\prime}}:=\langle\hat{c}_{k}\hat{c}_{k^{\prime}}^{\dagger}\rangle=\chi_{k}\delta_{kk^{\prime}}.$ Explicitly, $\chi_{k}$ is obtained as follows:
$\chi_{k}(t)=\begin{cases}1-e^{-\Gamma t/2}&(k<0),\\ 0&(k=0),\\ 1&(k>0).\end{cases}$ (25)
@@
The Gaussian form (22) of the density matrix turns out to be extremely powerful when later calculating the entanglement content. We emphasize that for bosons, the mixed state in Eq. (19) is not Gaussian anymore, despite having exactly the same form.
III.2 Entanglement measures